# A Light-Cone Approach to Higher-Order Cosmological Observables

### v1

@@ 1 Introduction
At first order in cosmological perturbation theory, the impact of Large-Scale Structure on the luminosity distance–redshift relation has been extensively studied. Pioneer works [4, 5, 6, 7, 8, 9] derived expressions for the weak-lensing convergence, Shapiro time-delay, Sachs-Wolfe, and Doppler contributions. These results have been used to forecast the magnitude of cosmic variance in distance measurements and its correlation function [10, 11]. First-order calculations have proven adequate for many applications, but as observational uncertainties shrink towards the percent level and beyond, second-order effects become non-negligible.
@@
With the aim of solving the above-mentioned issues, here we report a new method based on the GLC gauge to compute relativistic corrections to cosmological observables and, using it, we develop a comprehensive derivation of the luminosity distance–redshift relation at second order in cosmological perturbation theory and as seen by a free-falling observer using a gauge-invariant and divergences free approach. Building upon the GLC formalism, we systematically expand the distance–redshift relation to second order and isolate the observer dependent contributions, as was done in [37] at linear order.
Going into details, we will address three interrelated technical goals:
  • Identification of the observer-position terms. We will systematically derive all second-order metric and geodesic perturbations evaluated at the observer’s spacetime position. These include monopole and higher-multipole contributions to the redshift and luminosity distance, which have often been neglected or treated heuristically in previous works.
  • Second-order gauge fixing at the observer position. By building upon the general gauge transformation rules for cosmological perturbations, we implement a consistent gauge choice at the observer position to second order. This involves specifying residual gauge functions that uniquely fix time and spatial reparameterizations on the observer’s worldline, thereby eliminating unphysical degrees of freedom in the local expansion of the metric.
  • Gauge invariance and cancellation of divergences. We will demonstrate that, once observer-position terms and gauge conditions are properly accounted for (see the two points above), all would-be divergent contributions cancel in the final expression for the observed luminosity distance-redshift relation. In particular, we systematically show that all the terms at the observer position safely do not induce any divergences, ensuring a manifestly finite and gauge-invariant result. Notably, these cancellations occur without any assumption regarding the theory of gravity or the matter content of the Universe.
@@
This paper is organized as follows. In Sect. 2, we describe two possible approaches to cosmological perturbation theory: the standard one, where space-time is foliated into space-like hyper-surfaces, and the light-cone one, where space-time is foliated into light-like hyper-surfaces. After a proper splitting of the perturbations according to their symmetry properties, we compute the gauge transformations in both the approaches, working up to the second perturbative order. We also establish a fully non-linear dictionary between the standard and light-cone perturbations. In Sect. 3, we start by defining the GLC gauge at a fully non-linear order. Then, we impose the GLC gauge-fixing conditions at first and second order within the light-cone perturbation theory, ending up with the forms that gauge modes needed to transform from a generic gauge to the perturbative definition of the GLC one. The standard counterpart of such a gauge fixing, namely the OSG, is then defined in Sect. 4, where we point out its similarities and differences wrt the SG. In Sect. 5, we recall how the GLC coordinates turn out to be a very powerful tool to compute cosmological observables to any order in perturbation theory. In particular, we provide exact, non-perturbative expressions for the redshift and the angular distance. Then, we compute a set of first- and second-order gauge-invariant variables associated with the GLC gauge and use them to define a new approach to evaluate higher-order cosmological observables. As an application of this new perturbative scheme, in Sect. 6 we focus on the computation of the angular distance–redshift relation as seen by a free-falling observer. Thanks to a complete gauge-fixing procedure at the observer position, we show a definite way to solve the long-standing issue of divergences associated with the observer. To conclude, we provide the ultimate, gauge-invariant expression for the angular distance–redshift relation, whose source terms are systematically matched with the literature. Finally, the main conclusions and possible future directions of investigations are summarized in Sect. 7. In Appendix A, we prove that the observer defined by the OSG is free-falling, as well as the one of the SG.
@@ 2.1 Definition of perturbations
Here we are indicating with $D_{ij}$ the covariant derivative wrt the spatial background metric,
@@ 2.3 Map between standard and light-cone perturbation theory
To conclude this section, we establish a dictionary between the standard metric perturbations $f^{(n)}_{\mu\nu}$ and the light-cone ones $g^{(n)}_{\mu\nu}$. In order to do so, exploiting the usual diffeomorphism law
@@
$\begin{cases}a^{2}L=-2\big(\phi-\frac{1}{2}\mathcal{C}_{rr}-\mathcal{B}_{r}\big)\\ aM=-\mathcal{B}_{rr}-\mathcal{C}_{rr}\\ N=\mathcal{C}_{rr}\\ aV_{a}=-\mathcal{B}_{a}-\mathcal{C}_{ra}\\ U_{a}=\mathcal{C}_{ra}\\ \delta\gamma_{ab}=\mathcal{C}_{ab}\end{cases}\qquad\Rightarrow\qquad\begin{cases}\phi=-\frac{1}{2}(a^{2}L+N+2aM)\\ \mathcal{B}_{r}=-N-aM\\ \mathcal{C}_{rr}=N\\ \mathcal{B}_{a}=-U_{a}-aV_{a}\\ \mathcal{C}_{ra}=U_{a}\\ \mathcal{C}_{ab}=\delta\gamma_{ab}\end{cases}\,\,\,\,\,\,\,\,.$ (2.76)
@@ 3 The Geodesic Light-Cone Gauge
In this section, we will define the so-called Geodesic Light-Cone gauge at a fully non-linear level. Then, we will show the realizations of the linearized and quadratic versions of such a gauge fixing within the light-cone perturbation theory. In particular, we will provide the explicit expressions of the the gauge modes needed to fix the GLC gauge at first and second order.
@@ 3.1 Fully non-linear framework
$g_{\mu\nu}=\begin{pmatrix}0&-\Upsilon&\mathbf{0}\\ -\Upsilon&\Upsilon^{2}+\mathcal{U}^{2}&-\mathcal{U}_{b}\\ \mathbf{0}&-\mathcal{U}^{\text{T}}_{a}&g_{ab}\end{pmatrix}\qquad,\qquad g^{\mu\nu}=\begin{pmatrix}-1&-1/\Upsilon&-\mathcal{U}^{b}/\Upsilon\\ -1/\Upsilon&0&\mathbf{0}\\ -(\mathcal{U}^{a})^{\text{T}}/\Upsilon&\mathbf{0}&g^{ab}\end{pmatrix}$ (3.2)
@@ 3.2 Gauge modes at first order
The condition $V^{(1)}_{a}=0$ can be then be safely obtained by
@@
$\hat{\chi}_{(1)}$ $=-\int_{\tau}^{\tau_{\text{o}}}\text{d}\tau^{\prime}\,\hat{v}^{(1)}+\hat{\chi}^{(1)}_{0}$ (3.12)
@@ 3.3 Gauge modes at second order
Using Eq. (3.16) we can now impose the constraints $\tilde{v}^{(2)}=\hat{\tilde{v}}^{(2)}=0$ in Eqs. (2.67). From this, we obtain the following expressions for $\chi_{(2)}$ and $\hat{\chi}_{(2)}$:
@@ 4 The Observational Synchronous Gauge
In this section we will show how to impose the analogous of the GLC gauge in standard perturbation theory at second order. This fixing is called the Observational Synchronous Gauge, first introduced in [37] at first order in perturbation theory. We will then examine the conceptual similarities and differences between this gauge and the Synchronous Gauge (SG).
@@ 4.2 Second order
Let us note that $\phi^{(2)}\neq 0$, whereas this is not the case at first order. However, as first pointed out in [35], the proper time $\tau$ of the GLC gauge is non-perturbatively equal to the one of the SG. This implies that, at any perturbative order, the observers of the SG and OSG are both free-falling and then physically equivalent. Therefore, they measure the same physical observables. For the sake of completeness, and as a non-trivial check of our new perturbation theory, we will prove the above point that both the SG and OSG observers are free-falling in Appendix A by working at the perturbative level within light-cone perturbation theory.
@@
In order to show that, let us first derive the SG counterpart on the perturbed light-cone background. We have that the non-perturbative definition of the SG is given by $\phi^{(n)}=0$, $B^{(n)}=0$ and $B^{(n)}_{i}=0$ for any perturbative order $n$. As a consequence, using Eqs. (2.76), the non-perturbative version of the SG in the perturbative light-cone framework is given by
@@
Let us now make an brief remark and first look at the gauge transformations needed to fix the GLC gauge. Starting from Eqs. (2.76), and using the expression of $\phi^{(1)}$ in Eqs. (3.12), we obtain the first-order gauge mode to go to GLC gauge written in standard perturbation theory:
@@
This shows how $\xi^{\tau}_{(1)}$ corresponds to same time-shift of the SG, obtained by requiring $\tilde{\phi}^{(1)}=0$ and $\tilde{B}^{(1)}=0$ in Eqs. (2.22). This first result already show at first order how the GLC proper time and the SG one are identical.
In general, taking the limit $\tau\rightarrow\tau_{\text{o}}$ in Eqs. (3.12) at first order and in Eqs. (3.16), (3.17) and (3.20) at second order, we see that the only non-zero gauge modes are the time ones, modulo a non-zero integration constant in $\xi^{w}_{\text{o}}$ and $\xi^{a}_{\text{o}}$. Thus, a priori the only difference between a general gauge and the GLC one would be in the definition of the proper time at the observer. Then, as a matter of fact, with a little of algebra one can prove that the first and second-order time gauge modes needed to move from the SG to the GLC one (or the OSG) are zero. Indeed, using the first of Eqs. (4.6), we immediately see that $\xi^{\tau}_{(1)}=0$.
At second order, we can just use Eqs. (3.11) and (4.6) into (3.19) and (3.20), to show that, after some algebraic manipulations we get,
@@
which is consequence of the fact that the two gauges share the same proper time and are physically equivalent at the observer position, but differ at a generic time different from the one of the observer. At the second order, we also have non-zero expressions for $\xi^{w}_{(2)}$, $\chi_{(2)}$ and $\hat{\chi}_{(2)}$, that we do not show here for sake of brevity.
@@ 5 Gauge-Invariant Variables and Observables
In this section, we will first recall that cosmological observables admit non-perturbative expressions once the GLC gauge is fixed. As mentioned, this is ensured by the fact that, in the GLC gauge, observer and source are connected through $\tilde{\theta}^{a}=\text{constant}$ geodesics on a $w=\text{constant}$ past light-cone of a free-falling observer. We will then compute two sets of gauge-invariant variables which correspond to the first- and second-order versions of the GLC gauge. Finally, we will provide an unified treatment of all the GLC approaches to cosmological perturbations currently available in the literature.
@@ 5.1 Cosmological observables in the Geodesic Light-Cone gauge
in [36] it has been shown that the redshift in GLC coordinates is exactly given by
@@ 5.2 Gauge-invariant variables on the light-cone
In particular, starting from Eqs. (3.12), (3.16), (3.17) and (3.20), which give the dependence of the gauge fields $\xi^{\mu}_{(1)}$ and $\xi^{\mu}_{(2)}$ on the metric perturbations in order to move from a general gauge to the GLC gauge, we can perform a gauge transformation on the original perturbative light-cone variables using these fields. In this way, we obtain a set of variables which are automatically gauge invariant.
Proceeding as above, the set of gauge-invariant variables corresponding to the linearized version of the fully non-linear GLC gauge of Eq. (3.1) was already obtained in [37] with the following result:
@@
Evaluation in terms of Bardeen potentials.
@@
$\xi^{w}_{(2)}$ $=\int_{\tau}^{\tau_{\text{o}}}\,\frac{\text{d}\tau^{\prime}}{a}\bigg[-4\Psi^{\text{I}}_{(2)}-\bar{\gamma}^{ab}\partial_{a}Q\partial_{b}Q+4\Psi^{\text{I}}\partial_{w}Q-8\Psi^{\text{I}}\Psi^{\text{A}}-4(\Psi^{\text{I}})^{2}\bigg]+6\Psi^{\text{I}}_{\text{o}}P_{\text{o}}-2\Psi^{\text{I}}P$
$\quad+2P_{\text{o}}\partial_{w}P_{\text{o}}+(P_{\text{o}}-2Q)\partial_{w}Q-Q(\partial_{w}P_{\text{o}}-\partial_{w}Q)+\xi^{a}_{(1)}\partial_{a}Q+w^{(2)}_{0}(w,\tilde{\theta}^{a})\,$,
@@
Let us remark that Eqs. (5.13) and (5.18) are relevant outcomes of the general second-order perturbation theory developed so far. Indeed, any light-like cosmological observable can be first computed at a fully non-linear level using GLC coordinates. Then, it can be expanded up to the desired perturbative order within the framework of light-cone perturbation theory. After that, in order to have an expression valid in any gauge, we can use Eqs. (5.7) and  (5.8). Finally, thanks to the gauge modes of Eqs. (5.13) and (5.18) we can write the general relativistic corrections in terms of the gauge-invariant gravitational potentials.

### v2

@@ 1 Introduction
At first order in cosmological perturbation theory, the impact of Large-Scale Structure (LSS) on the luminosity distance–redshift relation has been extensively studied. Pioneer works [4, 5, 6, 7, 8, 9] derived expressions for the weak-lensing convergence, Shapiro time-delay, Sachs-Wolfe, and Doppler contributions. These results have been used to forecast the magnitude of cosmic variance in distance measurements and its correlation function [10, 11]. First-order calculations have proven adequate for many applications, but as observational uncertainties shrink towards the percent level and beyond, second-order effects become non-negligible.
@@
With the aim of solving the above-mentioned issues, here we report a new method based on the GLC gauge to compute relativistic corrections to cosmological observables and, using it, we develop a comprehensive derivation of the luminosity distance–redshift relation at second order in cosmological perturbation theory and as seen by a free-falling observer using a gauge-invariant and divergences-free approach. Building upon the GLC formalism, we systematically expand the distance–redshift relation to second order and isolate the observer dependent contributions, as was done in [37] at linear order.
Going into details, we obtain the new results summarized below, addressing also three interrelated technical goals as listed in the last three points:
  • A new method to obtain generic observables to non-linear order. By computing light-like observables on the past light-cone of a geodesic observer, the results are automatically expressed in term of the observed light-cone and observed angles. Therefore, only an expansion around the observed redshift is needed to write observables entirely in terms of observed quantities. This generalizes and extends to second order the first-order analysis described in [37].
  • General relativistic effects on the light-cone. The development of a higher-order gauge-invariant perturbation theory directly on the observed past light-cone allows for a better understanding of its intrinsic structure. This is because, once rewritten in terms of gauge-invariant gravitational potentials, the light-cone gauge modes (and their derivatives) encode well-defined linear and non-linear relativistic corrections.
  • Identification of the observer-position terms. We will systematically derive all second-order metric and geodesic perturbations evaluated at the observer’s space-time position. These include monopole and higher-multipole contributions to the redshift and luminosity distance, which have often been neglected or treated heuristically in previous works.
  • Second-order gauge fixing at the observer position. By building upon the general gauge-transformation rules for cosmological perturbations, we implement a consistent gauge choice at the observer position to second order. This involves specifying residual gauge functions that uniquely fix time and spatial reparameterizations on the observer’s worldline, thereby eliminating unphysical degrees of freedom in the local expansion of the metric.
  • Gauge invariance and cancellation of divergences. We will demonstrate that, once observer-position terms and gauge conditions are properly accounted for (see the last two points above), all would-be divergent contributions cancel in the final expression for the observed luminosity distance-redshift relation. In particular, we systematically show that all the terms at the observer position safely do not induce any divergences, ensuring a manifestly finite and gauge-invariant result. Notably, these cancellations occur without any assumption regarding the theory of gravity or the matter content of the Universe.
@@
This paper is organized as follows. In Sect. 2, we describe two possible approaches to cosmological perturbation theory: the standard one, where space-time is foliated into space-like hyper-surfaces, and the light-cone one, where space-time is foliated into light-like hyper-surfaces. After a proper splitting of the perturbations according to their symmetry properties, we compute the gauge transformations in both the approaches, working up to the second perturbative order. We also establish a fully non-linear dictionary between the standard and light-cone perturbations. In Sect. 3, we start by defining the GLC gauge at a fully non-linear order. Then, we impose the GLC gauge-fixing conditions at first and second order within the light-cone perturbation theory, ending up with the forms that gauge modes needed to transform from a generic gauge to the perturbative definition of the GLC one. The standard counterpart of such a gauge fixing, namely the OSG, is then defined in Sect. 4, where we point out its similarities and differences wrt the SG. In Sect. 5, we recall how the GLC coordinates turn out to be a very powerful tool to compute cosmological observables to any order in perturbation theory. In particular, we provide exact, non-perturbative expressions for the redshift and the angular distance. Then, we compute a set of first- and second-order gauge-invariant variables associated with the GLC gauge and use them to define a new approach to evaluate higher-order cosmological observables. As an application of this new perturbative scheme, in Sect. 6 we focus on the computation of the angular distance–redshift relation as seen by a free-falling observer. Thanks to a complete gauge-fixing procedure at the observer position, we show a definite way to solve the long-standing issue of divergences associated with the observer. To conclude, we provide the ultimate, gauge-invariant expression for the angular distance–redshift relation, whose source terms are systematically matched with the literature. Finally, the main conclusions and possible future directions of investigation are summarized in Sect. 7. In Appendix A, we prove that the observer defined by the OSG is free-falling, as well as the one of the SG. In Appendix B, for the sake of clarity, we provide a table summarizing the most relevant symbols and notations defined throughout the paper.
@@ 2.1 Definition of perturbations
Here we are indicating with $\nabla_{i}$ the covariant derivative wrt the spatial background metric, $D_{ij}$ is the traceless operator
@@ 2.3 Map between standard and light-cone perturbation theory
To conclude this section, we establish a dictionary between the standard metric perturbations $g^{(n)}_{\mu\nu}$ and the light-cone ones $f^{(n)}_{\mu\nu}$. In order to do so, exploiting the usual diffeomorphism law
@@
$\begin{cases}a^{2}L=-2\big(\phi-\frac{1}{2}\mathcal{C}_{rr}-\mathcal{B}_{r}\big)\\ aM=-\mathcal{B}_{r}-\mathcal{C}_{rr}\\ N=\mathcal{C}_{rr}\\ aV_{a}=-\mathcal{B}_{a}-\mathcal{C}_{ra}\\ U_{a}=\mathcal{C}_{ra}\\ \delta\gamma_{ab}=\mathcal{C}_{ab}\end{cases}\qquad\Rightarrow\qquad\begin{cases}\phi=-\frac{1}{2}(a^{2}L+N+2aM)\\ \mathcal{B}_{r}=-N-aM\\ \mathcal{C}_{rr}=N\\ \mathcal{B}_{a}=-U_{a}-aV_{a}\\ \mathcal{C}_{ra}=U_{a}\\ \mathcal{C}_{ab}=\delta\gamma_{ab}\end{cases}\,\,\,\,\,\,\,\,.$ (2.76)
@@ 3 The Geodesic Light-Cone Gauge
In this section, we will define the so-called Geodesic Light-Cone gauge at a fully non-linear level. Then, we will show the realizations of the linearized and quadratic versions of such a gauge fixing within the light-cone perturbation theory. In particular, we will provide the explicit expressions of the gauge modes needed to fix the GLC gauge at first and second order.
@@ 3.1 Fully non-linear framework
$g_{\mu\nu}=\begin{pmatrix}0&-\Upsilon&\mathbf{0}\\ -\Upsilon&\Upsilon^{2}+\mathcal{U}^{2}&-\mathcal{U}_{b}\\ \mathbf{0}^{\text{T}}&-\mathcal{U}^{\text{T}}_{a}&g_{ab}\end{pmatrix}\qquad,\qquad g^{\mu\nu}=\begin{pmatrix}-1&-1/\Upsilon&-\mathcal{U}^{b}/\Upsilon\\ -1/\Upsilon&0&\mathbf{0}\\ -(\mathcal{U}^{a})^{\text{T}}/\Upsilon&\mathbf{0}^{\text{T}}&g^{ab}\end{pmatrix}$ (3.2)
@@ 3.2 Gauge modes at first order
The condition $V^{(1)}_{a}=0$ can be then safely obtained by
@@
$\hat{\chi}_{(1)}$ $=-\int_{\tau}^{\tau_{\text{o}}}\text{d}\tau^{\prime}\,\hat{v}^{(1)}+\hat{\chi}^{(1)}_{0}\,$, (3.12)
@@ 3.3 Gauge modes at second order
Using Eq. (3.16), we can now impose the constraints $\tilde{v}^{(2)}=\hat{\tilde{v}}^{(2)}=0$ in Eqs. (2.67). From this, we obtain the following expressions for $\chi_{(2)}$ and $\hat{\chi}_{(2)}$:
@@ 4 The Observational Synchronous Gauge
In this section, we will show how to impose the analogous of the GLC gauge in standard perturbation theory at second order. This fixing is called the Observational Synchronous Gauge, first introduced in [37] at first order in perturbation theory. We will then examine the conceptual similarities and differences between this gauge and the Synchronous Gauge (SG).
@@ 4.2 Second order
Let us note that $\phi^{(2)}\neq 0$, whereas this is not the case at first order. However, as first pointed out in [35], the proper time $\tau$ of the GLC gauge is non-perturbatively equal to the one of the SG. This implies that, at any perturbative order, the observers of the SG and OSG are both free-falling and then physically equivalent. Therefore, they measure the same physical observables. For the sake of completeness, and as a non-trivial check of our new perturbation theory, we will prove the above point that both the SG and OSG observers are free-falling in Appendix A by working at the perturbative level within the light-cone perturbation theory.
@@
In order to show this, let us first derive the SG counterpart on the perturbed light-cone background. We have that the non-perturbative definition of the SG is given by $\phi^{(n)}=0$, $B^{(n)}=0$ and $B^{(n)}_{i}=0$ for any perturbative order $n$. As a consequence, using Eqs. (2.76), the non-perturbative version of the SG in the perturbative light-cone framework is given by
@@
Let us now make a brief remark and first look at the gauge transformations needed to fix the GLC gauge. Starting from Eqs. (2.76), and using the expression of $\phi^{(1)}$ in Eqs. (3.12), we obtain the first-order gauge mode to go to GLC gauge written in standard perturbation theory:
@@
This shows how $\xi^{\tau}_{(1)}$ corresponds to same time-shift of the SG, obtained by requiring $\tilde{\phi}^{(1)}=0$ and $\tilde{B}^{(1)}=0$ in Eqs. (2.22). This first result already shows at first order how the GLC proper time and the SG one are identical.
In general, taking the limit $\tau\rightarrow\tau_{\text{o}}$ in Eqs. (3.12) at first order and in Eqs. (3.16), (3.17) and (3.20) at second order, we see that the only non-zero gauge modes are the time ones, modulo a non-zero integration constant in $\xi^{w}_{\text{o}}$ and $\xi^{a}_{\text{o}}$. Thus, a priori the only difference between a general gauge and the GLC one would be in the definition of the proper time at the observer. Then, as a matter of fact, with a little of algebra one can prove that the first- and second-order time gauge modes needed to move from the SG to the GLC one (or the OSG) are zero. Indeed, using the first of Eqs. (4.6), we immediately see that $\xi^{\tau}_{(1)}=0$.
At second order, we can just use Eqs. (3.11) and (4.6) into (3.19) and (3.20), to show that, after some algebraic manipulations, we get
@@
which is consequence of the fact that the two gauges share the same proper time and are physically equivalent at the observer position, but differ at a generic time different from the one of the observer. At second order, we also have non-zero expressions for $\xi^{w}_{(2)}$, $\chi_{(2)}$ and $\hat{\chi}_{(2)}$, that we do not show here for the sake of brevity.
@@ 5 Gauge-Invariant Variables and Observables
In this section, we will first recall that cosmological observables admit non-perturbative expressions once the GLC gauge is fixed. As mentioned, this is ensured by the fact that, in the GLC gauge, observer and source are connected through $\tilde{\theta}^{a}=\text{constant}$ geodesics on a $w=\text{constant}$ past light-cone of a free-falling observer. We will then compute two sets of gauge-invariant variables which correspond to the first- and second-order versions of the GLC gauge. Finally, we will provide a unified treatment of all the GLC approaches to cosmological perturbations currently available in the literature.
@@ 5.1 Cosmological observables in the Geodesic Light-Cone gauge
in [36] it was shown that the redshift in GLC coordinates is exactly given by
@@ 5.2 Gauge-invariant variables on the light-cone
In particular, starting from Eqs. (3.12), (3.16), (3.17) and (3.20), which give the dependence of the gauge fields $\xi^{\mu}_{(1)}$ and $\xi^{\mu}_{(2)}$ on the metric perturbations in order to move from a general gauge to the GLC one, we can perform a gauge transformation on the original perturbative light-cone variables using these fields. In this way, we obtain a set of variables which are automatically gauge invariant.
Proceeding as above, the set of gauge-invariant variables corresponding to the linearized version of the fully non-linear GLC gauge of Eq. (3.1) was already obtained in [37] with the following results:
@@
Evaluation in terms of the Bardeen potentials.
@@
$\xi^{w}_{(2)}$ $=\int_{\tau}^{\tau_{\text{o}}}\,\frac{\text{d}\tau^{\prime}}{a}\bigg[-4\Psi^{\text{I}}_{(2)}-\bar{\gamma}^{ab}\partial_{a}Q\partial_{b}Q+4\Psi^{\text{I}}\partial_{w}Q-8\Psi^{\text{I}}\Psi^{\text{A}}-4(\Psi^{\text{I}})^{2}\bigg]+2\Psi^{\text{I}}_{\text{o}}P_{\text{o}}-2\Psi^{\text{I}}P$
$\quad+(P_{\text{o}}-2Q)\partial_{w}Q-Q(\partial_{w}P_{\text{o}}-\partial_{w}Q)+\xi^{a}_{(1)}\partial_{a}Q+w^{(2)}_{0}(w,\tilde{\theta}^{a})\,$,
@@
Let us remark that Eqs. (5.13) and (5.18) are relevant outcomes of the general second-order perturbation theory developed so far. Indeed, any light-like cosmological observable can be first computed at a fully non-linear level using GLC coordinates. Then, it can be expanded up to the desired perturbative order within the framework of light-cone perturbation theory. After that, in order to have an expression valid in any gauge, we can use Eqs. (5.7) and  (5.8). Finally, thanks to the gauge modes of Eqs. (5.13) and (5.18), we can write the general relativistic corrections in terms of the gauge-invariant gravitational potentials.